# Efficient Reasoning

### v1

@@ 1 Introduction
Tree-of-Thought (ToT) search strategies (Yao et al., 2023) enable Large Language Models to explore multiple reasoning paths beyond linear Chain-of-Thought prompting (Wei et al., 2022), improving performance on complex problems requiring backtracking or strategic lookahead. However, exploring vast reasoning trees remains computationally expensive, with many branches redundantly exploring semantically equivalent reasoning paths.
@@ 2 Related Work
Linear strategies such as Chain-of-Thought (CoT) (Wei et al., 2022) and Self-Consistent CoT (Wang et al., 2023) improved multi-step reasoning but explored limited paths. Tree-of-Thoughts (ToT) (Yao et al., 2023) expanded exploration using sampling, beam search, and MCTS (Cobbe et al., 2021; Hao et al., 2023; Wan et al., 2024), aligning with the "inference scaling law" (Hong et al., 2024; Lin et al., 2024; Fu et al., 2024; Cai et al., 2024). Dynamic Parallel Tree Search (DPTS) (Ding et al., 2025) improved parallel efficiency but still explored many redundant states. SSDP addresses this by pruning redundant paths before they are generated. Recent work has acknowledged that tree-based methods continue to suffer from computational overheads (Rufail et al., 2025), reinforcing the need for efficiency-focused optimizations like SSDP.
@@
Pruning improves reasoning even in linear CoT (Zhao et al., 2025). Methods like FETCH (Wang et al., 2025) merge semantically similar states but require fine-tuned models. SSDP provides a lightweight, general alternative, using only a small reward model and semantic merging to achieve large efficiency gains without task-specific tuning.
@@
Semantic Self-Consistency (Knappe et al., 2025) clusters completed reasoning chains post-hoc. In contrast, SSDP integrates semantic merging during tree growth, shaping the search process itself to make reasoning faster and more efficient.
@@ 3.1 Framework Overview
SSDP is designed to improve tree-based LLM reasoning efficiency by pruning semantically redundant nodes. It builds upon the parallel tree search framework of DPTS (Ding et al., 2025), inheriting dynamic expansion control.
@@
    Selection: The framework first selects up to $k$ promising leaf nodes from the search frontier to expand in parallel. This selection is guided by an Upper Confidence Bound (UCB) policy, defined by (Ding et al., 2025) as:
@@ 4 Experimental Setup
Our SSDP implementation builds on the official Dynamic Parallel Tree Search (DPTS) codebase (Ding et al., 2025), adding semantic merging during node expansion while retaining other architectural components for fair comparison.
@@ 4.1 Models, Datasets, and Baselines
Models. We evaluate four open-source, instruction-tuned models from the Llama and Qwen families: Llama-3.1-8B-Instruct (Team, 2024), Llama-3.2-3B-Instruct, Qwen2.5-1.5B-Instruct, and Qwen2.5-7B-Instruct (Bai et al., 2024). These were chosen to ensure our findings are robust across different model architectures and scales.
Datasets. Experiments were conducted on two standard multi-step reasoning benchmarks: GSM8K (Cobbe et al., 2021) and MATH500 (Hendrycks et al., 2021). These datasets have canonical scoring procedures, enabling direct comparison with prior work.
Baselines. We compare SSDP against four widely-used search methods to quantify its impact: (1) DPTS (Ding et al., 2025), the state-of-the-art parallel search framework; (2) MCTS, a standard Monte Carlo Tree Search implementation for LLM reasoning (Yao et al., 2023); (3) Best-of-N, which samples independent CoT traces and selects the best result (Cobbe et al., 2021); and (4) Beam Search, which maintains a fixed-size beam of the most promising reasoning paths (Yao et al., 2023).
@@ 5 Results
† Results were obtained from the DPTS paper (Ding et al., 2025).
@@ A.1 Reproducibility Statement
  • Prompts: All prompts were taken from the DPTS paper [11].

### v2

@@ 1 Introduction
Tree-of-Thought (ToT) search strategies (yao2023tree) enable Large Language Models to explore multiple reasoning paths beyond linear Chain-of-Thought prompting (wei2022chain), improving performance on complex problems requiring backtracking or strategic lookahead. However, exploring vast reasoning trees remains computationally expensive, with many branches redundantly exploring semantically equivalent reasoning paths.
@@ 2 Related Work
Linear strategies such as Chain-of-Thought (CoT) (wei2022chain) and Self-Consistent CoT (wang2022) improved multi-step reasoning but explored limited paths. Tree-of-Thoughts (ToT) (yao2023tree) expanded exploration using sampling, beam search, and MCTS (cobbe2021; hao2023; wan2024b), aligning with the "inference scaling law" (MLSYS2024_5321b1da; lin2024awqactivationawareweightquantization; fu2024breaksequentialdependencyllm; cai2024medusasimplellminference). Dynamic Parallel Tree Search (DPTS) (ding2025dynamicparalleltreesearch) improved parallel efficiency but still explored many redundant states. SSDP addresses this by pruning redundant paths before they are generated. Recent work has acknowledged that tree-based methods continue to suffer from computational overheads (rufail2025clearcontrastingtextualfeedback), reinforcing the need for efficiency-focused optimizations like SSDP.
@@
Pruning improves reasoning even in linear CoT (zhao2025pruningimprovereasoningrevisiting). Methods like FETCH (wang2025dontlosttreesstreamlining) merge semantically similar states but require fine-tuned models. SSDP provides a lightweight, general alternative, using only a small reward model and semantic merging to achieve large efficiency gains without task-specific tuning.
@@
Semantic Self-Consistency (knappe2025semanticselfconsistencyenhancinglanguage) clusters completed reasoning chains post-hoc. In contrast, SSDP integrates semantic merging during tree growth, shaping the search process itself to make reasoning faster and more efficient.
@@ 3.1 Framework Overview
SSDP is designed to improve tree-based LLM reasoning efficiency by pruning semantically redundant nodes. It builds upon the parallel tree search framework of DPTS (ding2025dynamicparalleltreesearch), inheriting dynamic expansion control.
@@
    Selection: The framework first selects up to $k$ promising leaf nodes from the search frontier to expand in parallel. This selection is guided by an Upper Confidence Bound (UCB) policy, defined by (ding2025dynamicparalleltreesearch) as:
@@ 4 Experimental Setup
Our SSDP implementation builds on the official Dynamic Parallel Tree Search (DPTS) codebase (ding2025dynamicparalleltreesearch), adding semantic merging during node expansion while retaining other architectural components for fair comparison.
@@ 4.1 Models, Datasets, and Baselines
Models. We evaluate four open-source, instruction-tuned models from the Llama and Qwen families: Llama-3.1-8B-Instruct (meta2024llama31), Llama-3.2-3B-Instruct, Qwen2.5-1.5B-Instruct, and Qwen2.5-7B-Instruct (qwen2024qwen2). These were chosen to ensure our findings are robust across different model architectures and scales.
Datasets. Experiments were conducted on two standard multi-step reasoning benchmarks: GSM8K (cobbe2021) and MATH500 (hendrycks2021math). These datasets have canonical scoring procedures, enabling direct comparison with prior work.
Baselines. We compare SSDP against four widely-used search methods to quantify its impact: (1) DPTS (ding2025dynamicparalleltreesearch), the state-of-the-art parallel search framework; (2) MCTS, a standard Monte Carlo Tree Search implementation for LLM reasoning (yao2023tree); (3) Best-of-N, which samples independent CoT traces and selects the best result (cobbe2021); and (4) Beam Search, which maintains a fixed-size beam of the most promising reasoning paths (yao2023tree).
@@ 5 Results
† Results were obtained from the DPTS paper (ding2025dynamicparalleltreesearch).
@@ A.1 Reproducibility Statement
  • Prompts: All prompts were taken from the DPTS paper [ding2025dynamicparalleltreesearch].